1.3 If a household’s income falls from R20 000 to R17 000 and its consumption falls from R18 000 to R15 000, then
its:
a) marginal propensity to consume is –0,67.
b) marginal propensity to consume is 0,88.
c) marginal propensity to consume is 0,20.
d) marginal propensity to save is zero.
e) marginal propensity to save is 0,12.
Answer #1

Option D.

  • The marginal propensity to save is the amount or the portion of income used up for saving rather than consumption.
  • It is calculated as the change in income divided by the change in the consumption. We then subtract the result obtained from 1.
  • MPS = (change in income ÷ change in consumption) - 1

= (18000 - 15000 ÷ 20,000 - 17,000 ) - 1

= ( 3000 ÷ 3000 ) - 1

= 1 - 1 = 0.

Hence the marginal propensity to save is equal to zero.